OpenStudy (anonymous):

Absolute value equations, need HELP. Solve for x: -3|2x + 6| = -12

OpenStudy (anonymous):

options: A. x=1 and x=5 B: X=-1 and x= -5 C: x=-9 and x =3 D: no solution

OpenStudy (anonymous):

Answer is A

OpenStudy (anonymous):

\[-3\left| 2x+6 \right|= -12\] \[2\left| x+3 \right|=\frac{ -12 }{ -3 } = 4\] \[\left| x+3 \right|=2\]\[\begin{matrix}x+3=2\rightarrow x=-1 \\ x+3=-2\rightarrow x=-5\end{matrix}\]